\begin{thebibliography}{10}\itemsep=-1pt

\bibitem{baltrusaitis2018openface}
Tadas Baltrusaitis, Amir Zadeh, Yao~Chong Lim, and Louis-Philippe Morency.
\newblock Openface 2.0: Facial behavior analysis toolkit.
\newblock In {\em 2018 13th IEEE International Conference on Automatic Face \&
  Gesture Recognition (FG 2018)}, pages 59--66. IEEE, 2018.

\bibitem{FACS}
P. Ekman.
\newblock What the face reveals: Basic and app studies of spon exp using the
  facial action coding system (facs).
\newblock {\em Ox Uni Press}, 1997.

\bibitem{fabiano2020gaze}
Diego Fabiano, Shaun Canavan, Heather Agazzi, Saurabh Hinduja, and Dmitry
  Goldgof.
\newblock Gaze-based classification of autism spectrum disorder.
\newblock {\em Pattern Recognition Letters}, 135:204--212, 2020.

\bibitem{hinduja2019fusion}
Saurabh Hinduja, Md~Taufeeq Uddin, Sk~Rahatul Jannat, Astha Sharma, and Shaun
  Canavan.
\newblock Fusion of hand-crafted and deep features for empathy prediction.
\newblock In {\em 2019 14th IEEE International Conference on Automatic Face \&
  Gesture Recognition (FG 2019)}, pages 1--4. IEEE, 2019.

\bibitem{kollias2022abaw}
Dimitrios Kollias.
\newblock Abaw: Valence-arousal estimation, expression recognition, action unit
  detection \& multi-task learning challenges.
\newblock {\em arXiv preprint arXiv:2202.10659}, 2022.

\bibitem{kollias2020analysing}
D Kollias, A Schulc, E Hajiyev, and S Zafeiriou.
\newblock Analysing affective behavior in the first abaw 2020 competition.
\newblock In {\em 2020 15th IEEE International Conference on Automatic Face and
  Gesture Recognition (FG 2020)(FG)}, pages 794--800.

\bibitem{kollias2019face}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Face behavior a la carte: Expressions, affect and action units in a
  single network.
\newblock {\em arXiv preprint arXiv:1910.11111}, 2019.

\bibitem{kollias2021distribution}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Distribution matching for heterogeneous multi-task learning: a
  large-scale face study.
\newblock {\em arXiv preprint arXiv:2105.03790}, 2021.

\bibitem{kollias2019deep}
Dimitrios Kollias, Panagiotis Tzirakis, Mihalis~A Nicolaou, Athanasios
  Papaioannou, Guoying Zhao, Bj{\"o}rn Schuller, Irene Kotsia, and Stefanos
  Zafeiriou.
\newblock Deep affect prediction in-the-wild: Aff-wild database and challenge,
  deep architectures, and beyond.
\newblock {\em International Journal of Computer Vision}, pages 1--23, 2019.

\bibitem{kollias2019expression}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Expression, affect, action unit recognition: Aff-wild2, multi-task
  learning and arcface.
\newblock {\em arXiv preprint arXiv:1910.04855}, 2019.

\bibitem{kollias2021affect}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Affect analysis in-the-wild: Valence-arousal, expressions, action
  units and a unified framework.
\newblock {\em arXiv preprint arXiv:2103.15792}, 2021.

\bibitem{kollias2021analysing}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Analysing affective behavior in the second abaw2 competition.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 3652--3660, 2021.

\bibitem{liao2000note}
Jason~JZ Liao and Jerry~W Lewis.
\newblock A note on concordance correlation coefficient.
\newblock {\em PDA journal of pharmaceutical science and technology},
  54(1):23--26, 2000.

\bibitem{srivastava2020recognizing}
Shivam Srivastava, Saandeep Aathreya~Sidhapur Lakshminarayan, Saurabh Hinduja,
  Sk~Rahatul Jannat, Hamza Elhamdadi, and Shaun Canavan.
\newblock Recognizing emotion in the wild using multimodal data.
\newblock In {\em Proceedings of the 2020 International Conference on
  Multimodal Interaction}, pages 849--857, 2020.

\bibitem{zafeiriou2017aff}
Stefanos Zafeiriou, Dimitrios Kollias, Mihalis~A Nicolaou, Athanasios
  Papaioannou, Guoying Zhao, and Irene Kotsia.
\newblock Aff-wild: Valence and arousal ‘in-the-wild’challenge.
\newblock In {\em Computer Vision and Pattern Recognition Workshops (CVPRW),
  2017 IEEE Conference on}, pages 1980--1987. IEEE, 2017.

\end{thebibliography}

\documentclass[10pt,twocolumn,letterpaper]{article}

\usepackage[pagenumbers]{cvpr} 
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}

\usepackage[pagebackref,breaklinks,colorlinks]{hyperref}

\usepackage[capitalize]{cleveref}
\crefname{section}{Sec.}{Secs.}
\Crefname{section}{Section}{Sections}
\Crefname{table}{Table}{Tables}
\crefname{table}{Tab.}{Tabs.}

\def\cvprPaperID{*****} 
\def\confName{CVPR}
\def\confYear{2022}

\begin{document}

\title{Random Forest Regression for Continuous Affect Using Facial Action Units}

\author{Saurabh Hinduja\\
University of Pittsburgh\\
4200 Fifth Ave, Pittsburgh, PA 15260\\
{\tt\small sah273@pitt.edu}
\and
Shaun Canavan\\
University of South Florida\\
4202 E Fowler Ave, Tampa, FL 33620\\
{\tt\small scanavan@usf.edu}
\and
Liza Jivnani\\
University of South Florida\\
10811, mc Kinley dr, Tampa, FL 33612\\
{\tt\small ljivnani@usf.edu}
\and
Sk Rahatul Jannat \\
University of South Florida\\
4202 E Fowler Ave, Tampa, FL 33620\\
{\tt\small jannat@usf.edu}
\and
V Sri Chakra Kumar\\
Cornell University\\
2 E Loop Rd, New York, NY 10044\\
{\tt\small vk386@cornell.edu}
}

\maketitle

\begin{abstract}
In this paper we describe our approach to the arousal and valence track of the 3rd Workshop and Competition on Affective Behavior Analysis in-the-wild (ABAW). We extracted facial features using OpenFace and used them to train a multiple output random forest regressor. Our approach performed comparable to the baseline approach on valence and outperformed in on arousal.
\end{abstract}


\section{Introduction}
\label{sec:intro}

It has been shown that use hand crafted features to train machine learning models perform significantly  well for affect recognition. Hinduja et al \cite{hinduja2019fusion} showed that fusing hand crafted features, for empathy prediction, worked well using multiple classifier types. Facial action units \cite{FACS} and gaze vectors have been shown to be very useful. Fabinao et al \cite{fabiano2020gaze} used gaze to classify the risk of autism. Srivastava et al \cite{srivastava2020recognizing} used facial action units and gaze vectors for recognizing affect in the wild. Motivated by these methods, we adopt a similar approach to the arousal and valence track of the 3rd Workshop and Competition on Affective Behavior Analysis in-the-wild (ABAW) \cite{kollias2022abaw, kollias2021analysing, kollias2020analysing, kollias2021distribution, kollias2021affect, kollias2019expression, kollias2019face, kollias2019deep, zafeiriou2017aff}. We propose to use hand-crafted features from the cropped frames in the dataset to train a random forest regressor to predict valence and arousal.
\section{Method}
\label{sec:Method}

Our proposed method for detecting frame level valence and arousal uses hand crafted features to train a multiple output random forest regressor. The hand crafted features we selected were action unit occurrence and intensity, the gaze vector, gaze angles, and head pose. To extract these features we used the publicly available Openface \cite{baltrusaitis2018openface}. We did not run open face on the raw videos but on the cropped frames provided. Due to the complexity and the volume of data it took us a long time to pre-process and prepare the hand crafted features for training.

Given the extracted features, we first created a vector of the hand crafted features $E = [l_x,l_y,l_z,r_x,r_y,r_z,avg_x,avg_y,t_x,t_y, t_z,o_x,o_y,o_z,AU_{I1}, \allowbreak AU_{I2},\dots, AU_{I17}, AU_{O1}, AU_{O2},\dots,AU_{O17}]$ where $l_{\{x,~y,~z\}}$ and $r_{\{x,~y,~z\}}$ are the (x, y, z) coordinates of the left and right eye gaze direction vector, respectively. $t_{\{x,~y,~z\}}$ and $o_{\{x,~y,~z\}}$ are the (x, y, z) coordinates of the head pose translation and orientation vectors, respectively. $AU_{Ii}$ and $AU_{Oi}$ are the AU intensity and AU occurrence respectively. Next, we use the this vector to train a 250 trees random forest with multiple output regressor. Using a multiple output regressor helps to use a method not suitable for multiple output to predict multiple outputs. It does by fitting one regressor per output. 
\section{Experiments and Results}
\label{sec:exp}
\begin{table*}
    \centering
        \caption{Comparison of Valence-Arousal Challenge on Validation Set }
 \begin{tabular}{|c|c|c|c|}
 \hline
 Method & CCC-Valence & CCC-Arousal & PVA \\
 \hline
 Baseline(ResNet50) \cite{kollias2022abaw} & 0.31  & 0.17  & 0.24 \\
 \hline
 Ours  & 0.26  & 0.19  & 0.225 \\
 \hline
 \end{tabular}%

    \label{tab:ValidationResults}
\end{table*}
To compare with the baseline, our proposed approach was evaluated using concordance correlation coefficient (CCC) \cite{liao2000note}. Through evaluating the validation set of the challenge data, using the proposed approach, we achieved a CCC score of 0.26 and 0.19 for valence and arousal, respectively. As can be seen in Table \ref{tab:ValidationResults}, while our methods does not beat the baseline for valence, it does outperform it for arousal. The average CC score across both arousal and valence was 0.225, which is comparable to the baseline approach.
\section{Conclusion}
\label{sec:Con}
In this paper, we have shown that hand-crafted features such as actions units, and gaze are a promising approach to predicting continuous affect (i.e., arousal and valence). The proposed approach was comparable to the baseline on valence and outperformed in on arousal. Future work will include the fusion of hand-crafted and deep features using more recent works such as visual transformers.

{\small
\bibliographystyle{ieee_fullname}
\bibliography{egbib}
}

\end{document}